OpenStudy (anonymous):

Solve the system of equations below by graphing both equations with a pencil and paper. What is the solution? y = 2x - 3 y = -x + 3 A. (0, -3) B. (1, -1) C. (3, 0) D. (2, 1)

OpenStudy (anonymous):

Ah, graphing, ok so have you tried anything?

OpenStudy (anonymous):

Its undefined.

OpenStudy (daniellelovee):

same thing as before

OpenStudy (anonymous):

its D

OpenStudy (daniellelovee):

you got it

OpenStudy (anonymous):

Well, I'm not sure how you got that, as your question requires you to graph it out, so why not just graph it out and see :)

OpenStudy (daniellelovee):

there are 2 ways to solve this one that requires graphing and other that requires substituting options

OpenStudy (nincompoop):

WHERE ARE YOUR PENCIL AND PAPER?

OpenStudy (anonymous):

Mhm, well when you have an equation in the form of y = mx+b to graph it, for example lets take \[y=-\frac{ 2 }{ 3 }x+4\] to graph such an equation, we look at our given values, so y tells us how far up, and x tells us how far along on the following axis. m in this equation represents the slope (how steep the line is) and b is the y - intercept (where the line crosses the y - axis). So look we have a slope of \[-\frac{ 2 }{ 3 } \implies \frac{ rise }{ run }\] the -2 tells us how far up or down it's going and run is how far to the left or right it is.  so you would get something around these lines :p